# TactileNet: Bridging the Accessibility Gap with AI-Generated Tactile Graphics for Individuals with Vision Impairment

### v1

@@ Abstract
Tactile graphics are essential for providing access to visual information for the 43 million people globally living with vision loss, as estimated by global prevalence data. However, traditional methods for creating these tactile graphics are labor-intensive and struggle to meet demand. We introduce TactileNet, the first comprehensive dataset and AI-driven framework for generating tactile graphics using text-to-image Stable Diffusion (SD) models. By integrating Low-Rank Adaptation (LoRA) and DreamBooth, our method fine-tunes SD models to produce high-fidelity, guideline-compliant tactile graphics while reducing computational costs. Evaluations involving tactile experts show that generated graphics achieve 92.86% adherence to tactile standards and 100% alignment with natural images in posture and features. Our framework also demonstrates scalability, generating 32,000 images (7,050 filtered for quality) across 66 classes, with prompt editing enabling customizable outputs (e.g., adding/removing details). Our work empowers designers to focus on refinement, significantly accelerating accessibility efforts. It underscores the transformative potential of AI for social good, offering a scalable solution to bridge the accessibility gap in education and beyond.
@@
1 Introduction
Ensuring visual accessibility for individuals with visual impairments is an increasingly critical challenge in the digital age. According to the International Agency for the Prevention of Blindness (IAPB), an estimated 1.1 billion people worldwide were living with vision loss in 2020, a number projected to increase by 55% by 2050 due to population growth and aging IAPB (2020). Among these, 43 million people are blind, and 295 million have moderate to severe visual impairment, underscoring the urgent need for inclusive solutions. Many individuals with vision loss face significant barriers in education and daily life, as learning materials and information systems remain heavily reliant on visual content. This growing disparity highlights the necessity for scalable, innovative approaches to improve accessibility and bridge the gap for those with visual impairments.
Tactile graphics, which convey visual information through textured surfaces, play a crucial role in bridging this accessibility gap. To be effective, they must adhere to Braille Authority of North America (BANA) guidelines BANA (2025) to ensure clarity and usability. However, traditional production methods remain labor-intensive, time-consuming, and difficult to scale. While modern advancements, such as graphic design software CorelDRAW (2025); AdobeIllustrator (2025); TactileView (2025) and electronic embossers IndexBraille (2025); ViewPlus (2025), have improved efficiency, they fall short in meeting the growing demand for high-quality tactile materials. Automated solutions, including AI-driven approaches and refreshable tactile displays, offer promise but face challenges related to high costs and limited training datasets Mukhiddinov and Kim (2021). Consequently, these technologies have yet to be widely integrated into educational systems.
To address these scalability challenges, we introduce TactileNet, the first comprehensive digital dataset designed to train adapters for AI models that generate tactile images from text prompts or combined text and natural image inputs. Given the absence of paired datasets, these models are initially trained using text-to-image Stable Diffusion (SD) models Ho et al. (2020); Rombach et al. (2021) and later evaluated under image-to-image translation scenarios to ensure practical usability. This step is crucial for preserving essential structural information while omitting extraneous details like color and complex textures, which may hinder tactile perception.
A key innovation of our approach is the integration of Low-Rank Adaptation (LoRA) Hu et al. (2021) and Dreambooth Ruiz et al. (2023), which enhance model efficiency and precision for high-fidelity tactile graphic generation. LoRA reduces computational costs by adapting low-rank parameters, while Dreambooth enables personalized training with minimal data. The model’s ability to transform RGB images into tactile formats is validated through human expert feedback, demonstrating its potential to improve tactile graphic quality and accessibility. By automating and streamlining the creation process, our method allows designers to focus on refinement and customization, significantly reducing time and effort.
The contributions of this paper are threefold:
@@ 1 Introduction
  • Developing a novel methodology leveraging SD, LoRA, and Dreambooth to automate tactile graphic creation and editing through text, aiding tactile graphic designers.
@@
2 Related Work
2.1 Tactile Graphics: Current Approaches and Limitations
Recent advancements in tactile graphics production have significantly improved accessibility for blind and visually impaired (BVI) individuals. A key focus has been on ensuring the quality of tactile graphics, with methods like those proposed by Gonzalez et al. (2019) assessing image complexity to determine suitability for conversion. However, these approaches often rely on manual or semi-automated processes, which are labor-intensive and difficult to scale.
The integration of AI into accessibility tools has introduced innovative solutions, such as touchscreen tablets with audio feedback Guinness et al. (2019) and tactile graphics finders Felipe and Guerra-Gómez (2020). These tools enhance the comprehension of visual information through alternative modalities, yet they remain limited in addressing the full pipeline of tactile graphic production. For instance, video description tools Yuksel et al. (2020a, b) provide valuable context but do not automate the creation of tactile graphics.
2.2 GenAI and Deep Learning in Accessibility
Generative AI (GenAI), particularly deep learning (DL), has shown promise in automating complex tasks across various domains Kirillov et al. (2023); Dastjerdi et al. (2024); Khan et al. (2024). However, a major challenge in developing AI models for tactile image generation is the scarcity of high-quality paired datasets. The systematic review by Mukhiddinov and Kim (2021) highlights the limited availability of large-scale datasets and the high cost of refreshable tactile displays, both of which hinder advancements in AI-driven tactile graphics generation.
Text-to-image and image-to-image models, such as Stable Diffusion (SD) Ho et al. (2020); Rombach et al. (2021), have demonstrated the ability to generate high-quality visual content from textual or visual inputs. However, their application in tactile graphics remains underexplored, primarily due to the lack of specialized datasets and the unique structural constraints of tactile representation.
2.3 Gaps in Existing Solutions and Our Contribution
While existing tools address specific aspects of BVI accessibility, they lack a comprehensive solution for scalable, high-quality tactile graphic production. Our work bridges this gap by introducing a curated dataset designed to train AI models for automating tactile graphic creation. Unlike previous approaches, our method leverages SD models, enhanced with Low-Rank Adaptation (LoRA) Hu et al. (2021) and Dreambooth Ruiz et al. (2023), to ensure efficiency and precision. This approach not only reduces the time and effort required by tactile graphic designers but also establishes a new benchmark for inclusivity and accessibility in the field.
3 Methodology
In this section, we introduce the preliminaries and outline our methodology for curating the TactileNet dataset and adapting Stable Diffusion (SD) models to generate tactile graphics.
3.1 Preliminaries
3.1.1 Denoising Diffusion Probabilistic Models (DDPMs)
A DDPM Sohl-Dickstein et al. (2015); Ho et al. (2020) consists of two Markov chains: a forward chain that gradually adds noise to data, transforming it into a Gaussian distribution, and a reverse chain that learns to denoise and reconstruct the data Koller and Friedman (2009).
@@ 3.1.1 Denoising Diffusion Probabilistic Models (DDPMs)
3.1.2 Text-to-Image SD Models
@@
Text-to-Image SD models extend DDPMs, using textual prompts to guide image generation and enable text-to-visual synthesis. These models represent a convergence of vision-language technologies, where the structured approach of DDPMs is adapted to understand and generate images based on textual cues Radford et al. (2021). This innovative application involves not only the transformation of noise into structured images but also the integration of linguistic elements to ensure that the generated visuals accurately reflect the described scenarios Yang et al. (2024); Kawar et al. (2023); Du et al. (2022). The progression from data to noise and back to refined images exemplifies the adaptive use of diffusion techniques in bridging textual and visual modalities.
3.1.3 Fine-Tuning SD Models
@@
Fine-tuning adapts pre-trained models to specific tasks, enhancing performance in applications requiring precision and detail. Our framework leverages two methods: Low-Rank Adaptation (LoRA) Hu et al. (2021) and Dreambooth Ruiz et al. (2023), fine-tuned using Kohya’s Trainer Linaqruf (2023) on text-image pairs to refine the model’s ability to interpret textual prompts. During inference, the system supports both text-to-image and image-to-image generation (e.g., via the SD web interface AUTOMATIC1111 (2025)), enabling tactile graphic synthesis from textual descriptions or reference images. This flexibility ensures practical usability, bridging modalities for tactile design.
3.1.4 Low-Rank Adaptation (LoRA)
@@
LoRA is a parameter-efficient fine-tuning method, ideal for resource-constrained scenarios or small datasets. LoRA updates only a subset of the model’s parameters, enabling adaptation to new tasks or improved performance on specific data without comprehensive re-training. Initially, a pre-trained model, which has been trained on a large dataset, serves as the basis for adaptation. LoRA specifically targets the weight matrices in the attention and feed-forward layers of the model layers that are critical for performance. Instead of altering the original weight matrices directly, LoRA introduces a pair of low-rank matrices $A$ and $B$ for each weight matrix $W$ that needs adaptation. For a weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA introduces low-rank matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll m,n$. The adapted weight matrix becomes $W+\Delta W$, where $\Delta W=AB$. The low-rank approach ensures that only a fraction of the original parameters are updated, thus significantly reducing the number of trainable parameters. The original parameters of the pre-trained model remain frozen, meaning they are not updated during the adaptation process. Only the parameters of the low-rank matrices $A$ and $B$ (and possibly biases) are trainable, which requires considerably fewer computational resources compared to full model re-training, while maintaining high accuracy.
3.1.5 DreamBooth
@@
DreamBooth fine-tunes SD models using a small set of subject images, enabling the generation of customized images with high fidelity. It involves selecting a unique identifier for a specific subject, which conditions the model to associate the subject’s detailed characteristics and context with this identifier. During the fine-tuning process, the model is trained on a relatively small dataset featuring images of the subject. This targeted training adjusts the model’s parameters, enabling it to generalize the appearance and context of the new subject across various scenarios not directly presented in the training data. This method effectively integrates the subject into the model’s latent space, allowing for the generation of new, high-quality images of the subject in diverse contexts and settings upon request.
Both LoRA and DreamBooth tailor generative models to specific applications. LoRA is ideal for situations requiring limited resources or minor adaptations, affecting only a minimal number of parameters. DreamBooth, although more computationally intensive, enables deep integration of new subjects into the model, facilitating the creation of customized, contextually precise images from text descriptions.
3.2 Dataset Creation
@@
3.2.1 Data Collection
We curated the TactileNet dataset by sourcing tactile images from online digital libraries, including Perkins College for the Blind, the Tactile Graphics Image Library of American Printing House (TGIL APH), the Provincial Resource Centre for the Visually Impaired (PRCVI), and BTactile Perkins (2025); APH (2024); PRCVI (2024); BTactile (2025). These libraries are known for their high-quality, expert-designed tactile images, which are widely used in educational settings for the visually impaired.
Given the limited number of available images for certain classes (e.g., the ”dog” class in TGIL APH yielded only four images, including a circus dog), we expanded our dataset by adding visually similar images from online social platforms such as Pinterest. To ensure these additional images met the high standards necessary for tactile interpretation, each image was carefully compared against the benchmarks set by the tactile images from the aforementioned libraries, considered as the gold standard in tactile design. This approach ensured consistency in quality and educational value across the dataset. Figure 2 illustrates the process from initial image sourcing to final tactile graphics compilation. Through this careful process, we compiled a dataset of 1,029 tactile images across 66 classes to support robust model training. Key statistics, including class distribution, are provided in Table 2.
@@ 3.2.1 Data Collection
Throughout the development of our dataset, we actively collaborated with industry partners specializing in accessibility solutions to ensure the reliability and real-world applicability of our data. These partners included tactile graphics designers and educators with expertise in visual impairments, who provided critical insights into the practical needs of end-users. Their feedback ensured that our dataset not only meets technical standards but also aligns with the requirements of real-world educational settings for the visually impaired. This collaboration underscores our commitment to creating socially impactful AI solutions that address accessibility challenges.
Figure 3: Comprehensive workflow of our framework, starting with fine-tuning (left), where TactileNet data (tactile images, prompts) refine the SD model. The process transitions to the generation phase (right), applying fine-tuned adapters atop the frozen SD model for text-to-image and image-to-image tactile graphic generation.
3.2.2 Text Prompts Generation for Tactile Graphics
To prepare the dataset for fine-tuning, we generated text prompts for each tactile image using ChatGPT OpenAI (2025) integrated with DALL-E, following a structured template mentioned below.
Create a tactile graphic of an [object], specifically designed for individuals with visual impairments. The graphic should feature raised, smooth lines to delineate the [pattern/features], against a simplistic background to ensure stark contrast.
Here the [object] is replaced by the class name (e.g cat). The [patterns/features] are replaced by specific features (e.g whiskers, eyes, paws etc) in the tactile graphic. These prompts are then thoroughly reviewed to ensure they strictly adhere to our predefined template and whether they are faithful to the input tactile graphic, thereby maintaining consistency and relevance throughout the dataset.
3.3 Models Development and Image Generation
3.3.1 Fine-Tuning Individual Models for Each Category
Our dataset, comprising text-tactile image pairs, was used to fine-tune 66 distinct models (one per class) using LoRA and DreamBooth (Figure 3 (left)). The “tactile” identifier ensured alignment with tactile features during training. Implemented via the Kohya Trainer Linaqruf (2023), this process adapts the Stable Diffusion-based Anything V3 model AnythingV3 (2024); Rombach et al. (2021), originally fine-tuned for an anime-style generation. We selected Anything V3 after comparing generation quality across SD v1 and v1.5. The Anything V3 fine-tuned adapter produced superior results when used atop the base SD v1.5 model during the generation phase, as detailed in the following subsection.
3.3.2 Generation of Tactile Images
Following fine-tuning, tactile graphics are generated using the fine-tuned SD v1.5 model, as depicted in Figure 3 (right). Our generation phase operates under two settings: (1) class-specific prompts alone and (2) prompts combined with natural images. The inclusion of natural images significantly improved output quality, with 30% of graphics meeting quality standards compared to 15% for text-only prompts. These results highlight the importance of multimodal inputs for tactile graphic synthesis, as further evaluated in Section 4.3.
4 Experimental Settings

### v2

@@ Abstract
Tactile graphics are essential for providing access to visual information for the 43 million people globally living with vision loss. Traditional methods for creating these graphics are labor-intensive and cannot meet growing demand. We introduce TactileNet, the first comprehensive dataset and AI-driven framework for generating embossing-ready 2D tactile templates using text-to-image Stable Diffusion (SD) models. By integrating Low-Rank Adaptation (LoRA) and DreamBooth, our method fine-tunes SD models to produce high-fidelity, guideline-compliant graphics while reducing computational costs. Quantitative evaluations with tactile experts show 92.86% adherence to accessibility standards. Structural fidelity analysis revealed near-human design similarity, with an SSIM of 0.538 between generated graphics and expert-designed tactile images. Notably, our method preserves object silhouettes better than human designs (SSIM = 0.259 vs. 0.215 for binary masks), addressing a key limitation of manual tactile abstraction. The framework scales to 32,000 images (7,050 high-quality) across 66 classes, with prompt editing enabling customizable outputs (e.g., adding or removing details). By automating the 2D template generation step—compatible with standard embossing workflows—TactileNet accelerates production while preserving design flexibility. This work demonstrates how AI can augment (not replace) human expertise to bridge the accessibility gap in education and beyond. Code, data, and models will be publicly released to foster further research.
I INTRODUCTION
Providing accessible alternatives to visual information for individuals with visual impairments is an urgent societal challenge in the digital age. According to the International Agency for the Prevention of Blindness (IAPB), an estimated 1.1 billion people worldwide were living with vision loss in 2020, a number projected to increase by 55% by 2050 due to population growth and aging [1]. Among these, 43 million people are blind, and 295 million have moderate to severe visual impairment, underscoring the urgent need for inclusive solutions. Many individuals with vision loss face significant barriers in education and daily life, as learning materials and information systems remain heavily reliant on visual content. This growing disparity highlights the necessity for scalable, innovative approaches to improve accessibility and bridge the gap for those with visual impairments.
Tactile graphics—raised, textured representations of visual content—serve as a vital medium for conveying graphical information through touch, adhering to strict guidelines (e.g., line thickness, spacing) set by the Braille Authority of North America (BANA) [2] to ensure clarity and usability. Figure 1 illustrates an example of such graphics, highlighting how both expert-designed and model-generated images must differentiate between near-side and far-side limbs to convey posture and depth—an essential consideration in tactile design for side-view animal depictions. Despite advancements in design tools (e.g., CorelDRAW, Adobe Illustrator, TactileView) and embossing technologies (e.g., ViewPlus and Index Braille), manual tactile graphic creation remains prohibitively labor-intensive, with simple images requiring 10–15 minutes and complex images taking many hours to produce [3, 4]. Existing AI-driven solutions [5] are hindered by the absence of large-scale, domain-specific datasets, forcing reliance on synthetic data that fails to capture tactile design nuances. Consequently, these technologies have yet to be widely integrated into educational systems.
To address these scalability challenges, we introduce TactileNet, the first comprehensive digital dataset designed to train adapters for AI models that generate tactile images from text prompts or combined text and natural image inputs. Given the absence of paired datasets, these models are initially trained using text-to-image Stable Diffusion (SD) models [6, 7] and later evaluated under image-to-image translation scenarios to ensure practical usability. This step is crucial for preserving essential structural information while omitting extraneous details like color and complex textures, which may hinder tactile perception.
Refer to caption Refer to caption
(a) Sourced tactile graphic (b) Generated tactile graphic
Figure 1: Example tactile graphics of a horse. Left: Expert-designed tactile graphic from existing libraries. Right: Tactile graphic generated by our fine-tuned model.
@@ I INTRODUCTION
Our framework uniquely integrates Low-Rank Adaptation (LoRA) [8] and DreamBooth [9], enabling parameter-efficient fine-tuning on TactileNet while preserving the base model’s generative diversity—critical for handling rare objects. LoRA reduces computational costs by adapting low-rank parameters, while Dreambooth enables personalized training with minimal data. The model’s ability to transform RGB images into tactile formats is validated through human expert feedback, demonstrating its potential to improve tactile graphic quality and accessibility. By automating and streamlining the creation process, our method allows designers to focus on refinement and customization, significantly reducing time and effort. The contributions of this paper are threefold:
@@
  • Developing a novel methodology leveraging SD, LoRA, and Dreambooth to automate tactile graphic generation and text-based editing, aiding tactile designers.
@@
II RELATED WORK
II-A Tactile Graphics: Current Approaches and Limitations
Recent advances in tactile graphic production have prioritized accessibility for blind and visually impaired (BVI) communities. Early work focused on manual design principles, such as [10]’s framework for assessing image complexity to determine tactile suitability. While foundational, these methods rely heavily on human expertise—a bottleneck for scalability. Emerging tools like tactile graphics finders [11] and audio-tactile tablets [12] enhance content discovery but fail to automate the end-to-end design pipeline. Similarly, video description systems [13] improve contextual awareness but lack tactile translation capabilities, leaving a critical gap in accessible content creation.
II-B GenAI and Deep Learning in Accessibility
Generative AI (GenAI) has advanced accessibility through applications like real-time captioning [14] and scene segmentation [15]. However, tactile graphic generation poses unique challenges, as noted in [5]’s systematic review: (1) scarcity of paired datasets (text/image → tactile graphics), and (2) high costs of refreshable tactile displays. Text-to-image models like Stable Diffusion (SD) [7] offer promise but require adaptation to tactile constraints (e.g., suppressing textures, emphasizing edges). Recent parameter-efficient tuning methods like Low-Rank Adaptation (LoRA) [8] address data scarcity but remain untested in tactile domains.
II-C Gaps and Contributions
Existing solutions address fragments of the tactile design workflow but lack holistic automation. Our work bridges three key gaps:
  • Dataset Scarcity: We introduce TactileNet, the first dataset curated for text/image-to-tactile translation.
  • Structural Adaptation: While SD generates photorealistic images [6], we retrain it to prioritize tactile usability (simplified shapes, high-contrast edges).
  • Customization: Prior tools focus on static outputs; our framework enables prompt-based edits (e.g., “Remove background clutter”).
This establishes a new benchmark for scalable, AI-assisted tactile design.
III METHODOLOGY
This section covers TactileNet dataset curation and SD model adaptation for tactile graphic generation.
III-A Preliminaries
III-A1 Denoising Diffusion Probabilistic Models (DDPMs)
A DDPM [16, 6] consists of two Markov chains: a forward chain that gradually adds noise to data, transforming it into a Gaussian distribution, and a reverse chain that learns to denoise and reconstruct the data [17].
@@ III-A1 Denoising Diffusion Probabilistic Models (DDPMs)
III-A2 Text-to-Image SD Models
@@
Text-to-Image SD models extend DDPMs, using textual prompts to guide image generation and enable text-to-visual synthesis. These models represent a convergence of vision-language technologies, where the structured approach of DDPMs is adapted to understand and generate images based on textual cues [18]. This involves not only transforming noise into structured images but also integrating linguistic elements to ensure that the generated visuals accurately reflect the described scenarios [19, 20, 21].
III-A3 Fine-Tuning SD Models
@@
Fine-tuning adapts pre-trained models to specific tasks, enhancing performance in applications requiring precision and detail. Our framework leverages two methods: LoRA [8] and Dreambooth [9], fine-tuned using Kohya’s Trainer [22] on text-image pairs to refine the model’s ability to interpret textual prompts. During inference, the system supports both text-to-image and image-to-image generation (e.g., via the SD web interface [23]), enabling tactile graphic synthesis from textual descriptions or reference images. This flexibility ensures practical usability.
III-A4 Low-Rank Adaptation (LoRA)
@@
LoRA is a parameter-efficient fine-tuning method, ideal for resource-constrained scenarios or small datasets. LoRA updates only a subset of the model’s parameters, enabling adaptation to new tasks without comprehensive re-training. It introduces a pair of low-rank matrices $A$ and $B$ for each weight matrix $W$. For a weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA introduces $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll m,n$. The adapted weight matrix becomes $W+\Delta W$, where $\Delta W=AB$. This significantly reduces the number of trainable parameters. The original model parameters remain frozen.
III-A5 DreamBooth
@@
DreamBooth fine-tunes SD models using a small set of subject images, enabling the generation of customized images with high fidelity. It involves selecting a unique identifier for a specific subject, training the model to associate the subject’s characteristics with this identifier. This adjusts the model’s parameters to generalize the appearance of the new subject across diverse scenarios.
III-B Dataset Creation
@@
III-B1 Data Collection
We curated the TactileNet dataset by sourcing tactile images from online digital libraries, including Perkins College for the Blind, the Tactile Graphics Image Library of American Printing House (TGIL APH), the Provincial Resource Centre for the Visually Impaired (PRCVI), and BTactile [24, 25, 26, 27]. These libraries are known for high-quality, expert-designed tactile images.
Given the limited number of available images for certain classes (e.g., only four ”dog” images in TGIL APH), we expanded the dataset by adding visually similar images from online platforms like Pinterest. Each image was carefully compared against gold-standard tactile designs to ensure consistency in quality and educational value. Figure 3 illustrates the data sourcing process. The final dataset comprises 1,029 tactile images across 66 classes. Table I provides class distribution statistics.
@@ III-B1 Data Collection
Throughout dataset development, we collaborated with industry partners specializing in accessibility solutions, including tactile graphics designers and educators. Their feedback ensured that our dataset met both technical standards and practical requirements for real-world educational use.
Figure 4: Comprehensive workflow of our framework, starting with fine-tuning (left), where TactileNet data (tactile images, text prompts) refine the SD model. The process transitions to the generation phase (right), applying fine-tuned adapters atop the frozen SD model for text-to-image and image-to-image tactile graphic generation.
III-B2 Text Prompts Generation for Tactile Graphics
We generated text prompts for each tactile image using ChatGPT [28] integrated with DALL-E, following a structured template:
Create a tactile graphic of an [object], specifically designed for individuals with visual impairments. The graphic should feature raised, smooth lines to delineate the [patterns], against a simplistic background.
The [object] was replaced by the class name (e.g., cat) and [patterns] by specific features (e.g., whiskers, eyes, paws). All prompts were reviewed to ensure adherence to the template and alignment with the tactile graphic.
III-C Models Development and Image Generation
III-C1 Fine-Tuning Individual Models for Each Category
We fine-tuned 66 distinct models (one per class) using LoRA and DreamBooth (Figure 4, left). The “tactile” identifier ensured alignment with tactile features during training. Implemented via the Kohya Trainer [22], this process adapted the Stable Diffusion-based Anything V3 model [29, 7]. Anything V3 was selected after comparing generation quality across SD v1 and v1.5. The fine-tuned adapter yielded superior results when used atop the base SD v1.5 model.
III-C2 Generation of Tactile Images
Following fine-tuning, tactile graphics were generated using the fine-tuned SD v1.5 model (Figure 4, right). Two generation settings were used: (1) class-specific text prompts alone and (2) prompts combined with natural images. Including natural images doubled the output quality, with 30% of graphics meeting standards versus 15% for text-only prompts (Section IV-D).
IV EXPERIMENTAL SETTINGS